### v1

# On the complexity of matrix Putinar’s Positivstellensatz

Abstract.
This paper studies the complexity of matrix Putinar’s Positivstellensatz on the semialgebraic set that is given by the polynomial matrix inequality. Under the archimedeanness, we prove a polynomial bound on degrees of terms appearing in the representation of matrix Putinar’s Positivstellensatz. Estimates on the exponent and constant are given. As a byproduct, a polynomial bound on the convergence rate of matrix sum-of-squares relaxations is obtained, which resolves an open question raised by Dinh and Pham. When the constraining set is unbounded (the archimedeanness fails), we also prove a similar bound for the matrix version of Putinar–Vasilescu’s Positivstellensatz by exploiting homogenization techniques.
Key words and phrases:
polynomial optimization, matrix optimization, Moment-SOS relaxations, Positivstellensatz
2020 Mathematics Subject Classification:
90C23,65K05, 90C22, 90C26
1. Introduction
Optimization with polynomial matrix inequalities has broad applications [6, 12, 15, 16, 17, 23, 42, 43, 52, 53]. A fundamental question to solve it globally is how to efficiently characterize polynomial matrices that are positive semidefinite on the semialgebraic set described by the polynomial matrix inequality. To be more specific, given an $\ell$$-\ell$ symmetric polynomial matrix $F(x)$ in $x:=(x_{1},\dots,x_{n})$ and a semialgebraic set $K$, how can we certify that $F\succeq 0$ on $K$? Here, the set $K$ is given as
(1.1) $K=\{x\in\mathbb{R}^{n}\mid G(x)\succeq 0\},$
where $G(x)$ is an $m$-by-$m$ symmetric polynomial matrix.
When $F(x)$ is a scalar polynomial (i.e., $\ell=1$) and $G(x)$ is a diagonal matrix whose diagonal entries are polynomials $g_{1}(x)$, $\dots$, $g_{m}(x)$, $K$ can be equivalently described as
(1.2) $K=\{x\in\mathbb{R}^{n}\mid g_{1}(x)\geq 0,\dots,g_{m}(x)\geq 0\}.$
Then, it reduces to the classical Positivstellensatz in real algebraic geometry and polynomial optimization. A polynomial $p$ is said to be a sum of squares (SOS) if $p=p_{1}^{2}+\dots+p_{t}^{2}$ for polynomials $p_{1},\dots,p_{t}$. Schmüdgen [44] proved that if $K$ is compact and $F>0$ on $K$, then there exist SOS polynomials $\sigma_{\alpha}$ such that
(1.3) $F=\sum\limits_{\alpha\in\{0,1\}^{m}}\sigma_{\alpha}g_{1}^{\alpha_{1}}\cdots g_{m}^{\alpha_{m}}.$
Under the archimedeanness (a condition slightly stronger than the compactness, referring to Subsection 2.2), Putinar [38] showed that if $F>0$ on $K$, then there exist SOS polynomials $\sigma_{0},\sigma_{1},\dots,\sigma_{m}$ such that
(1.4) $F=\sigma_{0}+\sigma_{1}g_{1}+\sigma_{m}g_{m},$
which is referred to as Putinar’s Positivstellensatz in the literature.
The complexities of the above scalar Positivstellensatze have been studied extensively. Suppose that $K\subseteq(-1,1)^{n}$ and the degree of $F$ is $d$. For the degree $d$, denote
$\mathbb{N}_{d}^{n}:=\left\{\alpha\in\mathbb{N}^{n}\mid\left.\alpha_{1}+\cdots+\alpha_{n}\leq d\right\}.\right.$
For a polynomial $p=\sum\limits_{\alpha\in\mathbb{N}_{d}^{n}}p_{\alpha}\frac{|\alpha|!}{\alpha_{1}!\cdots\alpha_{n}!}x^{\alpha}$, denote the norm
$\|p\|=\max\{p_{\alpha}:\alpha\in\mathbb{N}_{d}^{n}\}.$
It was shown in [46] that if $F\geq F_{\min}$ on $K$, then for
$k\geq cd^{2}\left(1+\left(d^{2}n^{d}\frac{\|F\|}{F_{\min}}\right)^{c}\right),$
there exist SOS polynomials $\sigma_{\alpha}$ with $\deg(\sigma_{\alpha}g_{1}^{\alpha_{1}}\cdots g_{m}^{\alpha_{m}})\leq k$ such that (1.3) holds. Here, $c$ is a positive constant depending on constraining polynomials $g_{1},\dots,g_{m}$. Based on the above bound, Nie and Schweighofer [35] proved that under the archimedeanness, there exists a constant $c>0$ depending on $g_{1},\dots,g_{m}$ such that if
(1.5) $k\geq c\exp\left(\left(d^{2}n^{d}\frac{\|F\|}{F_{\min}}\right)^{c}\right),$
then (1.4) holds for some SOS polynomials $\sigma_{1}$, $\dots$, $\sigma_{m}$ with $\deg(\sigma_{1}g_{1})\leq k$ , $\dots$, $\deg(\sigma_{m}g_{m})\leq k$. Recently, Baldi and Mourrain improved the exponent bound (1.5) to a polynomial bound. It was shown in [2] that the same result holds if
(1.6) $k\geq\gamma d^{3.5nL}\left(\frac{\|F\|_{\max}}{F_{\min}}\right)^{2.5nL}.$
In the above, $\gamma$ is a positive constant depending on $n$, $g_{1},\dots,g_{m}$, $L$ is the Lojasiewicz exponent associated with $g_{1},\dots,g_{m}$ and $\|F\|_{\max}$ is the max norm on the cube $[-1,1]^{n}$. The bound (1.6) has been further improved in [3] by removing the dependency of the exponent on the dimension $n$ and providing estimates on the constant $\gamma$. There exist better degree bounds when the set $K$ is specified, seeing [1, 4, 12, 48, 49, 50]. We also refer to [26] when $F,g_{1},\dots,g_{m}$ admit the correlative sparse sparsity.
Based on Putinar’s Positivstellensatz, Lasserre [28] introduced the Moment-SOS hierarchy for solving polynomial optimization and proved its asymptotic convergence under the archimedeanness. The above degree bounds give estimates on convergence rates of the Moment-SOS hierarchy directly. We also refer to [8, 19, 20, 21, 32, 33, 41] for the finite convergence of the Moment-SOS hierarchy under some regular assumptions.
When $F$, $G$ are general polynomial matrices, a matrix version of Putinar’s Positivstellensatz was provided in [25, 43]. Under the archimedeanness, it was proved that if $F\succ 0$ on $K$, then there exist polynomial matrices $P_{i}\in\mathbb{R}[x]^{m\times\ell}$ ($i=1,\dots,t$) and an SOS matrix $S$ such that
(1.7) $F=S+P_{1}^{T}GP_{1}+\cdots P_{t}^{T}GP_{t}.$
A polynomial matrix $S(x)$ is SOS if $S(x)=H(x)^{T}H(x)$ for some polynomial matrix $H(x)$. For the special case that $G$ is diagonal (i.e., $K$ is as in (1.2)), the exponential bound (1.5) was generalized to the matrix representation (1.7) in [14]. When $G$ is a general polynomial matrix, the complexity of matrix Putinar’s Positivstellensatz (1.7) is open, to the best of author’s knowledge. We also refer to [7, 24, 45] for other types of SOS representations for polynomial matrices.
Contributions
This paper studies the complexity of matrix Putinar’s Positivstellensatz. Throughout the paper, we assume that
$K\subseteq\{x\in\mathbb{R}^{n}\mid 1-\|x\|_{2}^{2}\geq 0\},$
unless it is specified. Denote the scaled simplex
(1.8) $\bar{\Delta}:=\left\{x\in\mathbb{R}^{n}\mid 1+x_{1}\geq 0,\ldots,1+x_{n}\geq 0,\sqrt{n}-x_{1}-\cdots-x_{n}\geq 0\right\}.$
Note that the set $K\subset\bar{\Delta}$ in our settings. The degree of the polynomial matrix $G$ is denoted by $d_{G}$, i.e.,
(1.9) $d_{G}=\max\{\deg(G_{ij}),~{}1\leq i,j\leq m\}.$
For a positive integer $m$, denote
(1.10) $\theta(m)=\sum\limits_{i=1}^{m}\left(\prod_{k=m+1-i}^{m}\frac{k(k+1)}{2}\right).$
Let $\|F\|_{B}$ be the Bernstein norm of $F$ on the scaled simplex $\bar{\Delta}$ (referring to Section 3 for details). The set of all polynomial matrices with representations (1.7) is the quadratic module of $\ell$$-\ell$ polynomial matrices generated by $G(x)$ in $\mathbb{R}[x]^{\ell\times\ell}$, for which we denote by $\mbox{QM}[G]^{\ell}$. The set $\mbox{QM}[G]^{\ell}$ is archimedean if there exists a scalar $r>0$ such that $(r-\|x\|_{2}^{2})I_{\ell}\in\mathrm{QM}[G]^{\ell}$.
The following is our main result. It generalizes the recent results of scalar Putinar’s Positivstellensatz as in [2, 3] to the matrix case and gives the first degree bound for matrix Putinar’s Positivstellensatz.
Theorem 1.1.
Let $K$ be as in (1.1) with $m\geq 2$. Suppose $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix of degree $d$, and $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$. If $F(x)\succeq F_{\min}\cdot I_{m}\succ 0$ on $K$, then for
(1.11) $k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{7\eta+3},$
there exists an $\ell$$-\ell$ SOS matrix $S$ and polynomial matrices $P_{i}\in\mathbb{R}[x]^{m\times\ell}$ $(i=1,\dots,t)$ with
$\deg(S)\leq k,~{}\deg(P_{i}^{T}GP_{i})\leq k~{}(i=1,\dots,t),$
such that
(1.12) $F=S+P_{1}^{T}GP_{1}+\cdots P_{t}^{T}GP_{t}.$
In the above, $\kappa,\eta$ are respectively Lojasiewicz constant and exponent as in (5.3), and $C>0$ is a constant independent of $F$, $G$.
Positivity certificate (1.12) can be applied to construct matrix SOS relaxations for solving polynomial matrix optimization [15, 43]. As a byproduct of Theorem 1.1, we can get a polynomial bound on the convergence rate of matrix SOS relaxations (see Theorem 5.5), which resolves an open question raised in [10].
Remark. The assumption that $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$ is not serious. Under the archimedeanness of $\mbox{QM}[G]^{\ell}$, i.e., $(r-\|x\|_{2}^{2})I_{\ell}\in\mathrm{QM}[G]^{\ell}$ for some $r>0$, we have that $(1-\|x\|_{2}^{2})I_{\ell}\in\mathrm{QM}[G(\sqrt{r}x)]^{\ell}$. Theorem 1.1 remains true with $F,G$ replaced by $F(\sqrt{r}x)$, $G(\sqrt{r}x)$.
When $K$ is unbounded, the Putinar type SOS representation (1.7) may not generally hold. For the scalar case (i.e., $F$ is a scalar polynomial and $K$ is described by $g_{1}(x)\geq 0,\dots,g_{m}(x)\geq 0$), it was shown in [39, 40] that if the highest degree homogeneous term of $F$ is positive definite in $\mathbb{R}^{n}$ and $F>0$ on $K$, then there exists a degree $k$ such that $(1+\|x\|_{2}^{2})^{k}F$ has a representation (1.4), which is referred to as Putinar-Vasilescu’s Positivstellensatz. A polynomial complexity was proved for this type of representation by Mai and Magron [31]. Recently, a matrix version of Putinar–Vasilescu’s Positivstellensatz was established in [11].
In the following, we prove a similar bound as in Theorem 1.1 for matrix Putinar–Vasilescu’s Positivstellensatz. Let $F^{hom}$ denote the highest degree homogeneous term of $F$. The homogenization of $F$ is denoted by $\widetilde{F}$, i.e., $\widetilde{F}(\tilde{x}):=x_{0}^{\deg(F)}F(x/x_{0})$ for $\tilde{x}:=(x_{0},x)$. The notation $\lambda_{\min}(A)$ denotes the smallest eigenvalue of the matrix $A$.
Theorem 1.2.
Let $K$ be as in (1.1) with $m\geq 2$. Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix of degree $d$. If $F(x)\succ 0$ on $K$ and $F^{hom}\succ 0$ on $\mathbb{R}^{n}$, then for
$k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}(\theta(m)+2)^{3}(n+1)^{2}(\lceil d_{G}/2\rceil)^{6}\kappa^{7}d^{14\eta}\left(\frac{\|\widetilde{F}\|_{B}}{\widetilde{F}_{\min}}\right)^{7\eta+3},$
there exists an $\ell$$-\ell$ SOS matrix $S$ and polynomial matrices $P_{i}\in\mathbb{R}[x]^{m\times\ell}$ $(i=1,\dots,t)$ with
$\deg(S)\leq 2k+d,~{}\deg(P_{i}^{T}GP_{i})\leq 2k+d~{}(i=1,\dots,t),$
such that
(1.13) $(1+\|x\|_{2}^{2})^{k}F=S+P_{1}^{T}GP_{1}+\cdots P_{t}^{T}GP_{t}.$
In the above, $\kappa$, $\eta$ are respectively Lojasiewicz constant and exponent as in (5.8), $C>0$ is a constant independent of $F$, $G$ and $\widetilde{F}_{\min}$ is the optimal value of the optimization
(1.14) $\left\{\begin{array}[]{rl}\min&\lambda_{\min}(\widetilde{F}(\tilde{x}))\\ \mathit{s.t.}&(x_{0})^{2\lceil d_{G}/2\rceil-d_{G}}\widetilde{G}(\tilde{x})\succeq 0,\\ &x_{0}^{2}+x^{T}x=1.\\ \end{array}\right.$
Outline of proofs
The proof of Theorem 1.1 involves three major steps:
Step 1. When $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix and $K$ is described by finitely many scalar polynomial inequalities (i.e., $K$ is as in (1.2) ), we establish a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensatz (see Theorem 3.3). It improves the exponential bound given by Helton and Nie [14]. The proof is provided in Section 3.
Step 2. When $K$ is described by a polynomial matrix inequality (i.e., $K$ is as in (1.1)), we show that $K$ can be equivalently described by finitely many scalar polynomial inequalities with exact estimates on the quantity and degrees, by exploiting a proposition due to Schmüdgen [45]. Additionally, these polynomials belong to the quadratic module generated by $G(x)$ in $\mathbb{R}[x]$. This is discussed in Section 4.
Step 3. Finally, we prove Theorem 1.1 by applying the main result in Step 1 (i.e, Theorem 3.3) to $F$ and equivalent scalar polynomial inequalities for $K$ obtained in Step 2. We refer to 5.
The proof of Theorem 1.2 relies on Theorem 1.1 and homogenization techniques.
The remaining of the paper is organized as follows. Section 2 presents some backgrounds. Section 5 provides the proofs of Theorems 1.1, 1.2 and studies the convergence rate of matrix SOS relaxations. Conclusions and discussions are drawn in Section 6.
2. Preliminaries
2.1. Notations
The symbol $\mathbb{N}$ (resp., $\mathbb{R}$) denotes the set of nonnegative integers (resp., real numbers). Let $\mathbb{R}[x]:=\mathbb{R}\left[x_{1},\ldots,x_{n}\right]$ be the ring of polynomials in $x:=\left(x_{1},\ldots,x_{n}\right)$ with real coefficients and let $\mathbb{R}[x]_{d}$ be the ring of polynomials with degrees $\leq d$. For $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}$ and a degree $d$, denote $|\alpha|\coloneqq\alpha_{1}+\cdots+\alpha_{n}$,
$\mathbb{N}_{d}^{n}:=\left\{\alpha\in\mathbb{N}^{n}\mid|\alpha|\leq d\right\}.$
For a polynomial $p,\deg(p)$ denotes its total degree. The degree of a polynomial matrix $P=(P_{ij}(x))\in\mathbb{R}[x]^{\ell_{1}\times\ell_{2}}$ is denoted by $\deg(P)$, i.e.,
$\deg(P):=\max\{\deg(P_{ij}(x)),~{}1\leq i\leq\ell_{1},1\leq j\leq\ell_{2}\}.$
For a real number $t,\lceil t\rceil$ denotes the smallest integer greater than or equal to $t$. For a matrix $A$, $A^{T}$ denotes its transpose and $\|A\|_{2}$ denotes the spectral norm of $A$. A symmetric matrix $X\succeq 0$ (resp., $X\succ 0$) if $X$ is positive semidefinite (resp., positive definite). For a smooth function $p$ in $x$, its gradient is denoted by $\nabla p$.
2.2. Some basics in real algebraic geometry
In this subsection, we review some basics in real algebraic geometry. We refer to [7, 15, 24, 25, 29, 34, 43, 45] for more details.
A polynomial matrix $S$ is said to be a sum of squares (SOS) if $S=P^{T}P$ for some polynomial matrix $P(x)$. The set of all $\ell$$-\ell$ SOS polynomial matrices in $x$ is denoted as $\Sigma[x]^{\ell}$. For a degree $k$, denote the truncation
$\Sigma[x]_{k}^{\ell}\,\coloneqq\,\left\{\sum_{i=1}^{t}P_{i}^{T}P_{i}\mid t\in\mathbb{N},~{}P_{i}\in\mathbb{R}[x]^{\ell\times\ell},~{}2\deg(P_{i})\leq k\right\}.$
For an $m$-by-$m$ symmetric polynomial matrix $G$, the quadratic module of $\ell$$-\ell$ polynomial matrices generated by $G(x)$ in $\mathbb{R}[x]^{\ell\times\ell}$ is
$\mbox{QM}[G]^{\ell}=\left\{S+\sum_{i=1}^{t}P_{i}^{T}GP_{i}\mid t\in\mathbb{N},~{}S\in\Sigma[x]^{\ell},~{}P_{i}\in\mathbb{R}[x]^{m\times\ell}\right\}.$
The $k$th degree truncation of $\mbox{QM}[G]^{\ell}$ is
(2.1) $\mbox{QM}[G]_{k}^{\ell}=\left\{S+\sum_{i=1}^{t}P_{i}^{T}GP_{i}\mid t\in\mathbb{N},~{}S\in\Sigma[x]^{\ell}_{k},~{}P_{i}\in\mathbb{R}[x]_{\lceil k-\deg(G)\rceil/2}^{m\times\ell}\right\}.$
Note that $\Sigma[x]^{\ell}\subseteq\mbox{QM}[G]^{\ell}$. For the case $\ell=1$, we simply denote that
$\Sigma[x]_{k}:=\Sigma[x]_{k}^{1},~{}\mbox{QM}[G]_{k}:=\mbox{QM}[G]_{k}^{1}.$
The quadratic module $\mathrm{QM}[G]^{\ell}$ is said to be archimedean if there exists $r>0$ such that
$\left(r-\|x\|_{2}^{2}\right)I_{\ell}\in\mathrm{QM}[G]^{\ell}.$
The polynomial matrix $G$ determines the semialgebraic set
$K:=\left\{x\in\mathbb{R}^{n}\mid G(x)\succeq 0\right\}.$
If $\mbox{QM}[G]^{\ell}$ is archimedean, then the set $K$ must be compact. However, when $K$ is compact, it is not necessarily that $\mbox{QM}[G]^{\ell}$ is archimedean. The following is matrix Putinar’s Positivstellensatz.
Theorem 2.1 ([25, 43]).
Suppose that $G$ is an $m$-by-$m$ symmetric polynomial matrix such that $\mbox{QM}[G]^{\ell}$ is archimedean. If the $\ell$$-\ell$ symmetric polynomial matrix $F\succ 0$ on $K$, then $F\in\mbox{QM}[G]^{\ell}$.
2.3. Some technical propositions
In this subsection, we present two useful propositions that are used in our proofs. The following is the classical Lojasiewicz inequality.
Corollary 2.2 ([5, 30]).
Let $A$ be a closed and bounded semialgebraic set and $f$ and $g$ two continuous semialgebraic functions from $A$ to $\mathbb{R}$ such that $f^{-1}(0)\subset g^{-1}(0)$. Then there exist two positive constants $\eta$, $\kappa$ such that:
$|g(x)|^{\eta}\leq\kappa|f(x)|,~{}\forall x\in A.$
The smallest constant $\eta$ is called the Lojasiewicz exponent and $\kappa$ is the corresponding Lojasiewicz constant.
The Markov inequality below can be used to estimate gradients of polynomials.
Theorem 2.3 ([3, 27]).
Let $\bar{\Delta}$ be the scaled simplex as in (1.8) and let $p\in\mathbb{R}[x]$ be a polynomial with degree $\leq d$. Then:
$\max_{x\in\bar{\Delta}}\|\nabla p(x)\|_{2}\leq\frac{2d(2d-1)}{\sqrt{n}+1}\max_{x\in\bar{\Delta}}|p(x)|.$
3. Degree bounds for scalar polynomial inequalities
In this section, we consider the case that $G(x)$ is a diagonal matrix with diagonal entries $g_{1},\dots,g_{m}\in\mathbb{R}[x]$. Then, the set $K$ as in (1.1) can be equivalently described as
(3.1) $K=\{x\in\mathbb{R}^{n}\mid g_{1}(x),\dots,\dots,g_{m}(x)\geq 0\}.$
If the symmetric polynomial matrix $F(x)\succ 0$ on $K$, we prove a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensatz (1.7). It improves the exponential bound shown by Helton and Nie [14]. The proof generalizes the techniques recently developed in [2, 3] to the matrix case.
Let $\Delta^{n}$ be the $n$-dimensional standard simplex, i.e.,
$\Delta^{n}=\{x\in\mathbb{R}^{n}\mid x_{1}+\dots+x_{n}=1,x_{1}\geq 0,\dots,x_{n}\geq 0\}.$
Denote the scaled simplex by
$\bar{\Delta}:=\left\{x\in\mathbb{R}^{n}\mid\sqrt{n}-x_{1}-\cdots-x_{n}\geq 0,1+x_{1}\geq 0,\ldots,1+x_{n}\geq 0\right\}.$
For a degree $t$, let $\{B_{\alpha}^{(t)}(x):\alpha\in\mathbb{N}^{n}_{t}\}$ be the Bernstein basis of degree $t$ on $\bar{\Delta}$, i.e.,
$B_{\alpha}^{(t)}(x)=\frac{t!}{\alpha_{1}!\cdots\alpha_{n}!}(n+\sqrt{n})^{-t}\left(\sqrt{n}-x_{1}-\cdots-x_{n}\right)^{t-|\alpha|}\left(1+x_{1}\right)^{\alpha_{1}}\cdots\left(1+x_{n}\right)^{\alpha_{n}},$
For a degree $t\geq d$, the $\ell\times\ell$ symmetric polynomial matrix $F$ of degree $d$ can be expressed in the Bernstein basis $\{B_{\alpha}^{(t)}(x):\alpha\in\mathbb{N}^{n}_{t}\}$ as
$F=\sum_{\alpha\in\mathbb{N}_{t}^{n}}F_{\alpha}^{(t)}B_{\alpha}^{(t)}(x),$
where $F_{\alpha}^{(t)}$ are $\ell$$-\ell$ symmetric matrices. The Bernstein norm of $F$ with respect to $\{B_{\alpha}^{(t)}(x):\alpha\in\mathbb{N}^{n}_{t}\}$ is defined as
$\|F\|_{B,t}=\max\limits_{\alpha\in\mathbb{N}^{n}_{t}}\{\|F_{\alpha}^{(t)}\|_{2}\}.$
For convenience of notations, we denote $\|F\|_{B,d}$ by $\|F\|_{B}$. Similarly, $F$ can also be expressed in the monomial basis $\{x^{\alpha}:\alpha\in\mathbb{N}^{n}_{d}\}$ as
$F=\sum_{\alpha\in\mathbb{N}_{d}^{n}}F_{\alpha}\frac{d!}{\alpha_{1}!\cdots\alpha_{n}!}x^{\alpha}.$
The spectral norm of $F$ is defined as
$\|F\|_{2}=\max\limits_{\alpha\in\mathbb{N}^{n}_{d}}\{\|F_{\alpha}\|_{2}\}.$
Suppose that $p\in\mathbb{R}[x]$ is a homogeneous polynomial with $\deg(p)=d$. The Polya’s Positivstellensatz [36, 37] states that if $p\geq p_{\min}>0$ on the standard simplex $\Delta^{n}$, then the polynomial $(x_{1}+\cdots+x_{n})^{k}p(x)$ has only positive coefficients if
$k>d(d-1)\frac{\|p\|}{2p_{\min}}-d.$
This result has been generalized to the matrix case by Scherer and Hol [43]. It was shown in [43, Theorem 3] that if $F(x)$ is an $\ell$$-\ell$ homogeneous symmetric polynomial matrix with $\deg(F)=d$ and $F(x)\succeq F_{\min}\cdot I_{m}\succ 0$ on $\triangle^{n}$, then for
$k>d(d-1)\frac{\|F\|_{2}}{2F_{\min}}-d,$
all coefficient matrices of $(x_{1}+\cdots+x_{n})^{k}F(x)$ in the monomial basis are positive definite.
In the following, we present a dehomogenized version of matrix Polya’s Positivstellensatz in terms of the Bernstein basis.
Lemma 3.1.
Suppose $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$. If $F(x)\succeq F_{\min}\cdot I_{m}\succ 0$ on $\bar{\Delta}$, then for
$k>d(d-1)\frac{\|F\|_{B}}{2F_{\min}}-d,$
all coefficient matrices of $F(x)$ in the Bernstein basis $\{B_{\alpha}^{(k+d)}(x):\alpha\in\mathbb{N}^{n}_{k+d}\}$ are positive definite.
Proof.
Suppose that $F$ can be expressed in the Bernstein basis $\{B_{\alpha}^{(d)}(x):\alpha\in\mathbb{N}^{n}_{d}\}$ as
$F=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}^{(d)}B_{\alpha}^{(d)}(x),$
where $F_{\alpha}^{(d)}$ are $\ell$$-\ell$ symmetric matrices. Denote
$\widehat{F}(y_{1},\dots,y_{n+1})=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}^{(d)}\frac{d!}{\alpha_{1}!\cdots\alpha_{n}!}y^{\alpha}y_{n+1}^{d-|\alpha|}.$
For each $(y_{1},\dots,y_{n+1})\in\Delta^{n+1}$, let
$x_{i}=(n+\sqrt{n})y_{i}-1~{}(i=1,\dots,n).$
Then we know that
$\begin{array}[]{ll}\sqrt{n}-\sum\limits_{i=1}^{n}x_{i}&=\sqrt{n}-\sum\limits_{i=1}^{n}((n+\sqrt{n})y_{i}-1)\\ &=\sqrt{n}+n-(n+\sqrt{n})\sum\limits_{i=1}^{n}y_{i}\\ &=(\sqrt{n}+n)(1-\sum\limits_{i=1}^{n}y_{i})\\ &\geq 0,\\ \end{array}$
which implies that $x:=(x_{1},\dots,x_{n})\in\bar{\Delta}$. By computations, we know that for $(y_{1},\dots,y_{n+1})\in\Delta^{n+1}$, the following holds
$\widehat{F}(y_{1},\dots,y_{n+1})=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}^{(d)}B_{\alpha}^{(d)}(x)=F(x)\succeq F_{\min}\cdot I_{m}.$
Hence, $\widehat{F}\succeq F_{\min}I_{m}$ on $\Delta^{n+1}$. It follows from Theorem 3 [43] that for $k>d(d-1)\frac{\|\widehat{F}\|_{2}}{2F_{\min}}-d$, all coefficient matrices of $(y_{1}+\cdots+y_{n+1})^{k}\widehat{F}(y_{1},\dots,y_{n+1})$ in the monomial basis are positive definite, i.e., there exist $P_{\alpha}\succ 0$ such that
$(y_{1}+\cdots+y_{n+1})^{k}\widehat{F}(y_{1},\dots,y_{n+1})=\sum_{\alpha\in\mathbb{N}_{k+d}^{n}}P_{\alpha}\frac{(k+d)!}{\alpha_{1}!\cdots\alpha_{n}!}y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}(y_{n+1})^{k+d-|\alpha|}.$
By substituting
$y_{i}=\frac{x_{i}+1}{n+\sqrt{n}}~{}(i=1,\dots,n),\,y_{n+1}=1-\frac{1}{n+\sqrt{n}}\left(\sum\limits_{i=1}^{n}x_{i}+n\right)$
into the above identity, we get that
$F(x_{1},\dots,x_{n})=\sum_{\alpha\in\mathbb{N}_{k+d}^{n}}P_{\alpha}B_{\alpha}^{(k+d)}(x).$
Since $\|\widehat{F}\|_{2}=\|F\|_{B}$, the conclusion follows. ∎∎
Lemma 3.2.
Suppose $P(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(P)=d_{P}$. Let $\xi\in\mathbb{R}^{\ell}$ with $\|\xi\|_{2}=1$, and denote
$P^{(\xi)}(x):=\xi^{T}P(x)\xi.$
Then, we have that
(3.2) $\|P^{(\xi)}\|_{B,d_{P}}\leq\|P\|_{B}.$
Furthermore, there exists $\xi_{0}\in\mathbb{R}^{\ell}$ with $\|\xi_{0}\|_{2}=1$ such that
(3.3) $\|P^{(\xi_{0})}\|_{B,d_{P}}=\|P\|_{B}.$
Proof.
Suppoose that $P$ can be expressed as
$P=\sum_{\alpha\in\mathbb{N}^{n}_{d_{P}}}P_{\alpha}^{(d_{P})}B_{\alpha}^{(d_{P})}(x),$
where $P_{\alpha}^{(d_{P})}$ are $\ell$$-\ell$ symmetric matrices. Note that
$\begin{array}[]{ll}\|P^{(\xi)}\|_{B,d_{P}}=\|\xi^{T}P(x)\xi\|_{B,d_{P}}&=\|\sum\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\xi^{T}P_{\alpha}^{(d_{P})}\xi B_{\alpha}^{(d_{P})}(x)\|_{B,d_{P}}\\ &=\max\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\{|\xi^{T}P_{\alpha}^{(d_{P})}\xi|\}\\ &\leq\max\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\{\|P_{\alpha}^{(d_{P})}\|_{2}\}=\|P\|_{B}.\end{array}$
Thus, we have
$\|P^{(\xi)}\|_{B,d_{P}}\leq\|P\|_{B}.$
Since there exists $\xi_{0}\in\mathbb{R}^{\ell}$ with $\|\xi_{0}\|_{2}=1$ such that
$\|P\|_{B}=\max\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\{|\xi_{0}^{T}P_{\alpha}^{(d_{P})}\xi_{0}|\},$
then we have that
$\|P^{(\xi_{0})}\|_{B,d_{P}}=\max\limits_{\alpha\in\mathbb{N}^{n}_{d_{P}}}\{|\xi_{0}^{T}P_{\alpha}^{(d_{P})}\xi_{0}|\}=\|P\|_{B}.$
Denote
(3.4) $\mathcal{G}(x)=-\min\left(\frac{g_{1}(x)}{\left\|g_{1}(x)\right\|_{B}},\ldots,\frac{g_{m}(x)}{\left\|g_{m}(x)\right\|_{B}},0\right).$
For $x\in\mathbb{R}^{n}$, let $d(x,K)$ denote the Euclidean distance to the set $K$, i.e.,
$d(x,K)=\min\{\|x-y\|_{2},\,\,y\in K\}.$
If $\mathcal{G}(x)=0$, then we have $x\in K$ and $d(x,K)=0$. By Corollary 2.2, there exist positive constants $\eta$, $\kappa$ such that the following Lojasiewicz inequality holds:
(3.5) $d(x,K)^{\eta}\leq\kappa\mathcal{G}(x),\quad\forall x\in\bar{\Delta}.$
Note that
$d_{G}=\max\{\deg(g_{1}),\dots,\deg(g_{m})\}.$
In the following, we prove a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensatz (1.7) when $K$ is given by finitely many scalar polynomial inequalities, which improves the exponential bound as in [14].
Theorem 3.3.
Let $G(x)$ be a diagonal matrix with diagonal entries $g_{1},\dots,g_{m}\in\mathbb{R}[x]$ and let $K$ be as in (3.1). Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$, and $1-\|x\|_{2}^{2}\in\mbox{QM}[G]$. If $F(x)\succeq F_{\min}\cdot I_{\ell}\succ 0$ on $K$, then $F\in\mbox{QM}[G]_{k}^{\ell}$ for
(3.6) $k\geq C\cdot{8}^{7\eta}m^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{7\eta+3}.$
In the above, $\kappa$, $\eta$ are respectively Lojasiewicz constant and exponent as in (3.5) and $C>0$ is a constant independent of $F$, $g_{1},\dots,g_{m}$.
Proof.
Without loss of generality, we assume that $\left\|g_{i}\right\|_{B}=1$ $(i=1,\dots,m)$. Otherwise, we can scale $g_{i}$ by $1/\|g_{i}\|_{B}$ and the bound will not change. Let
$\delta=\frac{1}{{8}^{\eta}\kappa d^{2\eta}}\left(\frac{F_{\min}}{\|F\|_{B}}\right)^{\eta},~{}v=\frac{\delta F_{\min}}{20m\|F\|_{B}},~{}\lambda=5\|F\|_{B}/\delta.$
For the above $\delta$, $v$, it follows from Proposition 3.2 [3] that there exist $h_{i}\in\mathbb{R}[x]$ such that for $x\in\bar{\Delta}$, we have
  • (i)
    if $g_{i}(x)\geq 0,\left|h_{i}(x)\right|\leq 2v$.
  • (ii)
    if $g_{i}(x)\leq-\delta,\left|h_{i}(x)\right|\geq\frac{1}{2}$.
  • (iii)
    $\left\|h_{i}\right\|_{B}\leq 1$.
  • (iv)
    $h_{i}\in\Sigma[x]_{k}$ with $k\leq 10^{5}nd_{G}^{2}\delta^{-2}v^{-1}$.
Let
$P=F-\lambda\sum_{i=1}^{m}h_{i}g_{i}I_{\ell}.$
In the following, we show that $P\succeq\frac{1}{4}F_{\min}\cdot I_{\ell}$ on $\bar{\Delta}$. Equivalently, for each $\xi\in\mathbb{R}^{\ell}$ with $\|\xi\|_{2}=1$, we prove that for every $x\in\bar{\Delta}$, the following holds
(3.7) $F^{(\xi)}(x)-\lambda\sum_{i=1}^{m}h_{i}(x)g_{i}(x)=\xi^{T}(F(x)-\lambda\sum_{i=1}^{m}h_{i}(x)g_{i}(x)I_{\ell})\xi\geq\frac{1}{4}F_{\min}.$
In the above, $F^{(\xi)}(x):=\xi^{T}F(x)\xi.$
Case I. Suppose that $F^{(\xi)}(x)\leq\frac{3F_{\min}}{4}$. Then we know that $x\notin K$. Let $y\in K$ be the vector such that $\|x-y\|_{2}=d(x,K)$. Since $F(y)\succeq F_{\min}\cdot I_{\ell}$, we have that
$\begin{array}[]{ll}\frac{F_{\min}}{4}\leq F^{(\xi)}(y)-F^{(\xi)}(x)&\leq\max\limits_{s\in\bar{\Delta}}\|\nabla F^{(\xi)}(s)\|_{2}\|x-y\|_{2}\\ &\leq\frac{2d(2d-1)}{\sqrt{n}+1}\|x-y\|_{2}\max\limits_{s\in\bar{\Delta}}|F^{(\xi)}(s)|,\\ \end{array}$
where the last inequality is from Theorem 2.3. Note that $\max_{x\in\bar{\Delta}}|F^{(\xi)}|\leq\|F^{(\xi)}\|_{B,d}$ (cf. [13]). It follows from Lemma 3.2 that
$\frac{F_{\min}}{4}\leq\frac{2d(2d-1)}{\sqrt{n}+1}\|F^{(\xi)}\|_{B,d}\|x-y\|_{2}\leq 2d^{2}\|F\|_{B}\|x-y\|_{2},$
By the Lojasiewicz inequality (3.5), we have that
$G(x)\geq\kappa^{-1}\|x-y\|_{2}^{\eta}\geq\frac{1}{{8}^{\eta}\kappa d^{2\eta}}(\frac{F_{\min}}{\|F\|_{B}})^{\eta}=\delta.$
Hence, there exists $i_{0}\in\{1,\dots,m\}$ such that $g_{i_{0}}(x)\leq-\delta$. By item (ii), we have $h_{i_{0}}(x)\geq\frac{1}{2}$. Since $\max_{x\in\bar{\Delta}}|F^{(\xi)}(x)|\geq F_{\min}$, we know that for $x\in\bar{\Delta}$,
$F^{(\xi)}(x)-\lambda\sum_{i=1}^{m}h_{i}(x)g_{i}(x)$ $\geq F^{(\xi)}(x)+\lambda\frac{1}{2}\delta-2\lambda v(m-1)$
$\geq F^{(\xi)}(x)+\lambda\frac{\delta}{4}+\lambda\left(\frac{\delta}{4}-2v(m-1)\right)$
$\geq-\|F^{(\xi)}(x)\|_{B,d}+\frac{5}{4}\|F^{(\xi)}(x)\|_{B,d}$
$\geq\frac{1}{4}\|F^{(\xi)}(x)\|_{B,d}\geq\frac{1}{4}\max_{x\in\bar{\Delta}}|F^{(\xi)}(x)|$
$\geq\frac{1}{4}F_{\min}.$
Case II. Suppose that $F^{(\xi)}(x)\geq\frac{3F_{\min}}{4}$. Then for every $x\in\bar{\Delta}$, the following holds
$F^{(\xi)}(x)-\lambda\sum_{i=1}^{m}h_{i}(x)g_{i}(x)\geq\frac{3}{4}F_{\min}-2m\lambda v\geq\frac{1}{4}F_{\min}.$
Thus, we know that $F^{(\xi)}(x)\geq\frac{1}{4}F_{\min}>0$ for all $x\in\bar{\Delta}$. Note that $\xi\in\mathbb{R}^{n}$ with $\|\xi\|_{2}=1$ is arbitrary. We have $P\succeq\frac{1}{4}F_{\min}\cdot I_{\ell}$ on $\bar{\Delta}$.
Since $\deg(h_{i})\leq 10^{5}nd_{G}^{2}\delta^{-2}v^{-1}$ $(i=1,\dots,m)$, we have that
$\begin{array}[]{ll}\deg(h_{i}g_{i})\leq 10^{5}nd_{G}^{3}\delta^{-2}v^{-1}&\leq 2\cdot 10^{6}mnd_{G}^{3}\delta^{-3}\frac{\|F\|_{B}}{F_{\min}}\\ &\leq 2\cdot 10^{6}\cdot{8}^{3\eta}mnd_{G}^{3}\kappa^{3}d^{6\eta}(\frac{\|F\|_{B}}{F_{\min}})^{3\eta+1}.\\ \end{array}$
By Lemma 3.2, there exists $\xi_{0}\in\mathbb{R}^{n}$ with $\|\xi_{0}\|_{2}=1$ such that $\|P\|_{B}=\|\xi_{0}^{T}P\xi_{0}\|_{B,\deg(P)}$. Then, we get that
$\begin{array}[]{ll}\|P\|_{B}&=\|\xi_{0}^{T}P\xi_{0}\|_{B,\deg(P)}\\ &=\|F^{(\xi_{0})}-\lambda\sum_{i=1}^{m}h_{i}g_{i}\|_{B,\deg(P)}\\ &\leq\|F^{(\xi_{0})}\|_{B,\deg(P)}+\lambda\sum_{i=1}^{m}\|h_{i}g_{i}\|_{B,\deg(P)}\\ &\leq\|F^{(\xi_{0})}\|_{B,\deg(F)}+\lambda\sum_{i=1}^{m}\|h_{i}\|_{B,\deg(P)-\deg(g_{i})}\|g_{i}\|_{B,\deg(g_{i})}\\ &\leq\|F\|_{B}+\lambda m\leq(1+5\delta^{-1}m)\|f\|_{B}\\ &\leq{8}^{\eta+1}m\kappa d^{2\eta}(\frac{\|F\|_{B}}{F_{\min}})^{\eta}\|F\|_{B}.\\ \end{array}$
It follows from Lemma 3.1 that for
$\begin{array}[]{ll}k&\geq(16\cdot 10^{6})^{2}\cdot{8}^{7\eta}m^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}(\frac{\|F\|_{B}}{F_{\min}})^{7\eta+3}\\ &>\left(2\cdot 10^{6}\cdot{8}^{3\eta}mnd_{G}^{3}\kappa^{3}d^{6\eta}(\frac{\|F\|_{B}}{F_{\min}})^{3\eta+1}\right)^{2}{8}^{\eta+2}m\kappa d^{2\eta}(\frac{\|F\|_{B}}{F_{\min}})^{\eta}\frac{\|F\|_{B}}{F_{\min}}\\ &\geq\deg(P)^{2}\frac{\|P\|_{B}}{\frac{1}{2}F_{\min}},\\ \end{array}$
all coefficient matrices of $P(x)$ in the Bernstein basis $\{B_{\alpha}^{(k)}(x):\alpha\in\mathbb{N}_{k}^{n}\}$ are positive definite, i.e., there exist matrices $P_{\alpha}^{(k)}\succ 0$ such that
$F-\lambda\sum_{i=1}^{m}h_{i}g_{i}I_{\ell}=\sum_{\alpha\in\mathbb{N}_{k}^{n}}P_{\alpha}^{(k)}B_{\alpha}^{k}(x).$
Since $1+x_{1},\cdots,1+x_{n},\sqrt{n}-x_{1}-\cdots-x_{n}\in\mbox{QM}[1-\|x\|_{2}^{2}]$ and $1-\|x\|_{2}^{2}\in\mbox{QM}[G]$, the conclusion holds for sufficiently large $C$. ∎∎
Remark. If the linear independence constraint qualification holds at every $x\in K$, it was shown in [2, 3] that the Lojasiewicz exponent as in (3.5) is equal to 1, i.e., $\eta=1$. Then, under the assumptions of Theorem 3.3, we know that $f\in\mbox{QM}[G]_{k}^{\ell}$ if
(3.8) $k\geq C\cdot m^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14}\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{10}.$
4. A scalar representation for polymomial matrix inequalities
In this section, we provide exact estimates on the quantity and degree bound for polynomials $d_{1},\dots,d_{t}$ with $d_{1},\dots,d_{t}\in\mbox{QM}[G]$ such that the polynomial matrix inequality $G(x)\succeq 0$ can be equivalently described by scalar polynomial inequalities $d_{1}\geq 0$, $\dots$, $d_{t}\geq 0$. This is a quantitative version of the following proposition due to Schmüdgen [45].
Proposition 4.1 (Proposition 9, [45]).
Let $G(x)$ be an $m$-by-$m$ symmetric polynomial matrix and $G(x)\neq 0$. Then there exist $m$-by-$m$ diagonal matrices $D_{r}$, $m$-by-$m$ symmetric matrices $X_{\pm,r}$ and polynomials $z_{r}\in$ $\Sigma[x],r=1,\ldots,t$, such that:
  • (i)
    $X_{+r}X_{-r}=X_{-r}X_{+r}=z_{r}I,D_{l}=X_{-r}GX_{-r}^{T},z_{r}^{2}G=X_{+r}D_{r}X_{+r}^{T}$,
  • (ii)
    For $x\in\mathbb{R}^{n},A(x)\geq 0$ if and only if $D_{r}(x)\geq 0$ for all $r=1,\ldots,t$.
Let
(4.1) $K=\{x\in\mathbb{R}^{n}\mid G(x)\succeq 0\}.$
A direct corollary of Proposition 4.1 is that there exist finite many polynomials in $\mbox{QM}[G]$ such that $K$ can be equivalently described by these polynomials. Denote
$\theta(m)=\sum\limits_{i=1}^{m}\left(\prod_{k=m+1-i}^{m}\frac{k(k+1)}{2}\right).$
Recall that
$d_{G}=\max\{\deg(G_{ij}),~{}1\leq i,j\leq m\}.$
In the following, we provide a quantitative version of Proposition 4.1.
Theorem 4.2.
Suppose $G(x)$ is an $m$-by-$m$ symmetric polynomial matrix with $m\geq 2$. Then there exist $d_{1},d_{2},\dots,d_{\theta(m)}\in QM[G]_{3^{m-1}d_{G}}$ such that
$K=\{d_{1}\geq 0,\dots,d_{\theta(m)}\geq 0\}.$
Proof.
We prove the result by applying the induction on the matrix size $m$. When $m=2$, the inequality $G(x)=[G_{ij}(x)]_{i,j=1,2}\succeq 0$ can be equivalently described as
$G_{11}\geq 0,~{}G_{22}\geq 0,~{}G_{11}(G_{11}G_{22}-G_{12}^{2})\geq 0,$
$G_{11}+2G_{12}+G_{22}\geq 0,~{}G_{22}(G_{11}G_{22}-G_{12}^{2})\geq 0,$
$(G_{11}+2G_{12}+G_{22})((G_{11}+2G_{12}+G_{22})G_{22}-(G_{12}+G_{22})^{2})\geq 0.$
We can verify that the above polynomials all belong to $\mbox{QM}[G]_{3d_{G}}$. Note that $\theta(2)=6$. Hence, it holds for $m=2$.
Suppose it is true for $m=k$ and consider $m=k+1$. For $i\in\{1,\ldots,n\}$, denote by $P_{i}$ the permutation matrix which permutes the first row and the $i$-th row. For $11\leq i<j\leq k+$ 1, denote the row operation matrix $Q^{(ij)}$ as
$(Q^{(ij)})_{uv}=\left\{\begin{array}[]{ll}1,&\text{if~{}}(u,v)=(1,1),\dots,(k+1,k+1)\text{~{}and~{}}(i,j),\\ 0,&\text{otherwise}.\\ \end{array}\right.$
Let
$T_{ii}=P_{i}~{}(i=1,\dots,k+1),~{}T_{ij}=P_{i}\cdot Q^{(ij)}~{}(1\leq i<j\leq k+1).$
One can check that for $11\leq i\leq j\leq k+$ 1, the above matrices $T_{ij}$ are all invertible and it holds that
$T_{ij}GT_{ij}^{\mathrm{T}}=\left[\begin{array}[]{cc}s_{ij}&(\beta^{(ij)})^{\mathrm{T}}\\ \beta^{(ij)}&H^{(ij)}\end{array}\right],$
where $H^{(ij)}$is a $k$-by-$k$ symmetric polynomial matrix, $\beta^{(ij)}\in\mathbb{R}[x]^{k}$ and
$s_{ii}=G_{ii}~{}(i=1,\dots,k+1),$
$s_{ij}=G_{ii}+G_{jj}+2G_{ij}~{}(1\leq i<j\leq k+1).$
Denote matrices
$X_{+}^{(ij)}=\left[\begin{array}[]{cc}s_{ij}&~{}0\\ \beta^{(ij)}&~{}s_{ij}\cdot I_{k}\end{array}\right],~{}X_{-}^{(ij)}=\left[\begin{array}[]{cc}s_{ij}&~{}0\\ -\beta^{(ij)}&~{}s_{ij}\cdot I_{k}\end{array}\right].$
By computations, we have that
(4.2) $X_{-}^{(ij)}X_{+}^{(ij)}=X_{+}^{(ij)}X_{-}^{(ij)}=s_{ij}^{2}\cdot I_{k+1},$
(4.3) $(X_{-}^{(ij)}T_{ij})\cdot G\cdot(X_{-}^{(ij)}T_{ij})^{\mathrm{T}}=G^{(ij)},$
(4.4) $s_{ij}^{4}T_{ij}GT_{ij}^{\mathrm{T}}=X_{+}^{(ij)}\cdot G^{(ij)}(X_{+}^{(ij)})^{\mathrm{T}},$
where
$G^{(ij)}=\left[\begin{array}[]{cc}s_{ij}^{3}&0\\ 0&s_{ij}\cdot\left(s_{ij}H^{(ij)}-\beta^{(ij)}(\beta^{(ij)})^{\mathrm{T}}\right)\end{array}\right].$
Let
$B^{(ij)}=s_{ij}\left(s_{ij}H^{(ij)}-\beta^{(ij)}(\beta^{(ij)})^{\mathrm{T}}\right)~{}(1\leq i\leq j\leq k+1).$
Since $T_{ij}$ are scalar matrices, we have
$\deg(s_{ij})\leq d_{G},~{}\deg(B^{(ij)})\leq 3d_{G}.$
In the following, we show that
(4.5) $K=\{x\in\mathbb{R}^{n}\mid s_{ij}(x)\geq 0,B^{(ij)}(x)\succeq 0,~{}1\leq i\leq j\leq k+1\}.$
If $x\in K$, i.e., $G(x)\succeq 0$, it follows from (4.3) that $G^{(ij)}(x)\succeq 0$, i.e.,
(4.6) $s_{ij}(x)\geq 0,B^{(ij)}(x)\succeq 0,\quad\forall 1\leq i\leq j\leq k+1.$
For the converse part, suppose that $x\in\mathbb{R}^{n}$ satisfies (4.6). If $s_{ij}=0$ for all $11\leq i\leq j\leq k+$ 1, then we have that $G_{ij}(x)=0$ for all $11\leq i,j\leq k+$ 1, and thus $G(x)=0$. If there exist $11\leq i_{0}\leq j_{0}\leq k+$ 1 such that $s_{i_{0}j_{0}}(x)\neq 0$, then the relation (4.4) gives that
(4.7) $G(x)=s_{i_{0}j_{0}}^{-4}(x)T_{ij}^{-1}X_{+}^{(i_{0}j_{0})}\cdot G^{(i_{0}j_{0})}(X_{+}^{(i_{0}j_{0})})^{\mathrm{T}}(T_{i_{0}j_{0}}^{\mathrm{T}})^{-1}.$
Since $G^{(i_{0}j_{0})}(x)\succeq 0$, we know that $G(x)\succeq 0$.
By the induction, each $B^{(ij)}$ is a $k$-by-$k$ symmetric polynomial matrix with $\deg(B^{(ij)})\leq 3^{k-1}d_{G}$ for $11\leq i\leq j\leq k+$ 1 and there exist $d_{1}^{(ij)},\dots,d_{\theta(k)}^{(ij)}\in QM[G]_{3^{k}d_{G}}$ such that
$\{x\in\mathbb{R}^{n}\mid B^{(ij)}(x)\succeq 0\}=\{x\in\mathbb{R}^{n}\mid d_{1}^{(ij)}(x)\geq 0,\dots,d_{\theta(k)}^{(ij)}\geq 0\}.$
Combining with the relation (4.5), we have that
$K=\{x\in\mathbb{R}^{n}\mid s_{ij}(x)\geq 0,d_{t}^{(ij)}(x)\succeq 0,~{}1\leq i\leq j\leq k+1,1\leq t\leq\theta(k)\}.$
The total number of the above constraints is
$\frac{(k+1)(k+2)}{2}+\frac{(k+1)(k+2)}{2}\theta(k),$
which equals to $\theta(k+1)$. Hence, we complete the proof.
Remark. There exist alternative ways to obtain finitely many scalar polynomial inequalities to describe $K$. For instance, consider the characteristic polynomial $det[\lambda I_{m}-G(x)]$ of $G(x)$. It can be expressed as
$det[\lambda I_{m}-G(x)]=\lambda^{m}+\sum\limits_{i=1}^{m}(-1)^{i}g_{i}(x)\lambda^{m-i}.$
Since $G(x)$ is symmetric, all eigenvalues are real and $det[\lambda I_{m}-G(x)]$ has only real roots and we know that $G(x)\succeq 0$ if and only if $g_{i}(x)\geq 0$ $(i=1,\dots,m)$. Hence, if polynomials $g_{2}(x),\dots,g_{m}(x)\in\mbox{QM}[G]$, the quantity $\theta(m)$ and the degree $3^{m-1}d_{G}$ as in Theorem 4.2 can be improved to $m$ and $md_{G}$, respectively. However, $g_{2}(x),\dots,g_{m}(x)$ typically do not belong to $\mbox{QM}[G]$.
5. Degree bounds for matrix Positivstellensatze
This section is mainly to prove Theorems 1.1 and 1.2. As a byproduct, we can get a polynomial bound on the convergence rate of matrix SOS relaxations for solving polynomial matrix optimization, which resolves an open question raised by Dinh and Pham [10]. Let
(5.1) $K=\{x\in\mathbb{R}^{n}\mid G(x)\succeq 0\},$
where $G(x)$ is an $m$-by-$m$ symmetric polynomial matrix.
5.1. Proof of Theorem 1.1
For the polynomial matrix $G(x)$, let $d_{1},\dots,d_{\theta(m)}$ be polynomials as in Theorem 4.2. Note that we assume that
$K\subseteq\{x\in\mathbb{R}^{n}\mid 1-\|x\|_{2}^{2}\geq 0\}.$
Denote
(5.2) $\mathcal{G}(x)=-\min\left(\frac{d_{1}(x)}{\left\|d_{1}(x)\right\|_{B}},\ldots,\frac{d_{\theta(m)}(x)}{\left\|d_{\theta(m)}(x)\right\|_{B}},0\right).$
If $\mathcal{G}(x)=0$, we have that $x\in K$, i.e., $d(x,K)=0$. By Corollary 2.2, there exist positive constants $\eta$, $\kappa$ such that the following Lojasiewicz inequality holds:
(5.3) $d(x,K)^{\eta}\leq\kappa\mathcal{G}(x),~{}\forall x\in\bar{\Delta}.$
In the following, we prove Theorem 1.1.
Theorem 5.1.
Let $K$ be as in (5.1) with $m\geq 2$. Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$, and $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$. If $F(x)\succeq F_{\min}\cdot I_{m}\succ 0$ on $K$, then $F\in\mbox{QM}[G]_{k}^{\ell}$ for
(5.4) $k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{7\eta+3}.$
In the above, $\kappa,\eta$ are respectively Lojasiewicz constant and exponent as in (5.3) and $C>0$ is a constant independent of $F$, $G$.
Proof.
By Theorem 4.2, we know that $d_{1},\dots,d_{\theta(m)}\in\mbox{QM}[G]_{3^{m-1}d_{G}}$ and
$K=\{d_{1}(x)\geq 0,\dots,d_{\theta(m)}(x)\geq 0\}.$
It follows from Theorem 3.3 that for
$k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{7\eta+3},$
there exist matrices $P_{\alpha}^{(k)}\succ 0$, SOS polynomial matrices $S_{i}\in\Sigma[x]_{k}^{\ell}$ ($i=0,\dots,\theta(m)+1$) with $\deg(S_{0})\leq k$, $\deg(d_{i}S_{i})\leq k~{}(i=1,\dots,\theta(m))$ such that
$F=S_{0}+d_{1}S_{1}+\cdots d_{\theta(m)}S_{\theta(m)}+\sum_{\alpha\in\mathbb{N}_{k}^{n}}P_{\alpha}^{(k)}B_{\alpha}^{k}(x).$
Since $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$, we know that $1-\|x\|_{2}^{2}\in\mbox{QM}[G]$. Note that $d_{1},d_{2},\dots,d_{\theta(m)}\in QM[G]$ and $1+x_{1},\cdots,1+x_{n},\sqrt{n}-x_{1}-\cdots-x_{n}\in\mbox{QM}[1-\|x\|_{2}^{2}]$. The proof is completed. ∎∎
Remark. When $G(x)$ is a diagonal polynomial matrix with diagonal entries $g_{1},\dots,g_{m}\in\mathbb{R}[x]$, i.e.,
$K=\{x\in\mathbb{R}^{n}\mid g_{1}(x)\geq 0,\dots,g_{m}(x)\geq 0\},$
Theorem 5.1 implies that for $k$ greater than the quantity as in (5.4), we have $F\in\mbox{QM}[G]_{k}^{\ell}$. However, the bound (3.6) provided in Theorem 3.3 is typically smaller than that as in (5.4), and the proof of Theorem 5.1 builds up on Theorem 3.3.
5.2. Proof of Theorem 1.2
For a polynomial matrix $F=\sum\limits_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}x^{\alpha}$ with $\deg(F)=d$, $F^{hom}$ denotes its highest-degree homogeneous terms, i.e.,
$F^{hom}(x)=\sum\limits_{\alpha\in\mathbb{N}^{n},|\alpha|=d}F_{\alpha}x^{\alpha}.$
The homogenization of $F$ is denoted by $\widetilde{F}$, i.e.,
$\widetilde{F}(\tilde{x})=\sum_{\alpha\in\mathbb{N}^{n}_{d}}F_{\alpha}x^{\alpha}x_{0}^{d-|\alpha|}$
for $\tilde{x}:=(x_{0},x)$. Let
(5.5) $\begin{array}[]{rcl}\widetilde{K}:=&\left\{\tilde{x}\in\mathbb{R}^{n+1}\left|\begin{array}[]{l}(x_{0})^{d_{0}}\widetilde{G}(\tilde{x})\succeq 0,\\ x_{0}^{2}+x^{T}x=1\end{array}\right.\right\},\\ \end{array}$
where $d_{0}=2\lceil d_{G}/2\rceil-d_{G}$. Consider the optimization
(5.6) $\left\{\begin{array}[]{rl}\min&\lambda_{\min}(\widetilde{F}(\tilde{x}))\\ \mathit{s.t.}&(x_{0})^{d_{0}}\widetilde{G}(\tilde{x})\succeq 0,\\ &x_{0}^{2}+x^{T}x=1,\\ \end{array}\right.$
where $\lambda_{\min}(\widetilde{F}(\tilde{x}))$ is the smallest eigenvalue of $\widetilde{F}(\tilde{x})$. Let $\widetilde{F}_{\min}$ denote the optimal value of (5.6).
Lemma 5.2.
Let $K$ be as in (5.1). Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$. If $F\succ 0$ on $K$ and $F^{hom}\succ 0$ on $\mathbb{R}^{n}$, then $\widetilde{F}_{\min}>0$ and $\widetilde{F}\succeq\widetilde{F}_{\min}\cdot I_{\ell}$ on $\widetilde{K}$.
Proof.
Suppose $\tilde{x}=(x_{0},x)\in\widetilde{K}$. If $x_{0}=0$, we have that
$\widetilde{F}(\tilde{x})=F^{hom}(x)\succ 0.$
If $x_{0}\neq 0$, then
$G(x/x_{0})=(x_{0})^{-d_{G}-d_{0}}(x_{0})^{d_{0}}\tilde{G}(\tilde{x})\succeq 0,$
which implies that $x/x_{0}\in K$. Since $F^{hom}\succ 0$ on $\mathbb{R}^{n}$, the degree of $F$ must be even and we know that
$\widetilde{F}(\tilde{x})=x_{0}^{d}F(x/x_{0})\succ 0.$
Since $\widetilde{K}$ is compact, the optimal value of (5.6) is positive, i.e., $\widetilde{F}_{\min}>0$. ∎∎
For the polynomial matrix $(x_{0})^{d_{0}}\widetilde{G}(\tilde{x})$, let $d_{1}(\tilde{x}),\dots,d_{\theta(m)}(\tilde{x})$ be the scalar polynomials as in Theorem 4.2. Denote
(5.7) $\mathcal{G}(\tilde{x})=-\min\left(\frac{d_{1}(\tilde{x})}{\left\|d_{1}(\tilde{x})\right\|_{B}},\ldots,\frac{d_{\theta(m)}(\tilde{x})}{\left\|d_{\theta(m)}(\tilde{x})\right\|_{B}},\frac{1-\|\tilde{x}\|_{2}^{2}}{\|1-\|\tilde{x}\|_{2}^{2}\|_{B}},\frac{\|\tilde{x}\|_{2}^{2}-1}{\|1-\|\tilde{x}\|_{2}^{2}\|_{B}},0\right).$
Note that $\mathcal{G}(\tilde{x})=0$ implies that $\tilde{x}\in\widetilde{K}$. By Corollary 2.2, there exist positive constants $\eta$, $\kappa$ such that for all $\tilde{x}$ satisfying $\sqrt{n+1}-x_{0}-x_{1}-\cdots-x_{n}\geq 0,1+x_{0}\geq 0,1+x_{1}\geq 0,\ldots,1+x_{n}\geq 0$, it holds that
(5.8) $d(\tilde{x},\widetilde{K})^{\eta}\leq\kappa\mathcal{G}(\tilde{x}).$
In the following, we prove Theorem 1.2. It provides a polynomial bound on the complexity of matrix Putinar–Vasilescu’s Positivstellensatz.
Theorem 5.3.
Let $K$ be as in (5.1) with $m\geq 2$. Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$. If $F\succ 0$ on $K$ and $F^{hom}\succ 0$ on $\mathbb{R}^{n}$, then $(1+\|x\|_{2}^{2})^{k}F\in\mbox{QM}[G]_{2k+d}^{\ell}$ for
(5.9) $k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}(\theta(m)+2)^{3}(n+1)^{2}(\lceil d_{G}/2\rceil)^{6}\kappa^{7}d^{14\eta}\left(\frac{\|\widetilde{F}\|_{B}}{\widetilde{F}_{\min}}\right)^{7\eta+3}.$
In the above, $\kappa$, $\eta$ are respectively Lojasiewicz constant and exponent as in (5.8), $C>0$ is a constant independent of $F$, $G$ and $\widetilde{F}_{\min}$ is the optimal value of (5.6).
Proof.
By Lemma 5.2, we have that $\widetilde{F}\succeq\widetilde{F}_{\min}\cdot I_{\ell}\succ 0$ on $\widetilde{K}$. Note that $\mbox{QM}[\|\tilde{x}\|^{2}-1]^{\ell}$ is archimedean and $K$ can be equivalently described as
$d_{1}(\tilde{x})\geq 0,\dots,d_{\theta(m)}(\tilde{x})\geq 0,1-\|\tilde{x}\|_{2}^{2}\geq 0,\|\tilde{x}\|_{2}^{2}-1\geq 0,$
By following the proof of Theorem 1.1 for $\widetilde{F}(x)$, we know that for $k$ satisfying (5.9), there exist $S\in\Sigma[\tilde{x}]^{\ell}$, $H\in\mathbb{R}[\tilde{x}]^{\ell\times\ell}$, $P_{i}\in\mathbb{R}[\tilde{x}]^{m\times\ell}$ ($i=1,\dots,t$) with
$\deg(S)\leq k,\,\deg(H)\leq k-2,\,\deg(P_{i}^{T}(x_{0})^{d_{0}}\widetilde{G}P_{i})\leq k~{}(i=1,\dots,t),$
such that
$\widetilde{F}(\tilde{x})=S(\tilde{x})+(\|\tilde{x}\|_{2}^{2}-1)\cdot H(\tilde{x})+\sum\limits_{i=1}^{t}P_{i}(\tilde{x})^{T}(x_{0})^{d_{0}}\widetilde{G}(\tilde{x})P_{i}(\tilde{x}).$
Replacing $\tilde{x}$ by $(1,x)/\sqrt{1+\|x\|_{2}^{2}}$ in the above identity, we get that
(5.10) $\begin{array}[]{ll}\frac{F(x)}{\sqrt{1+\|x\|_{2}^{2}}^{d}}=&\sum\limits_{i=1}^{t}P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)^{T}\frac{G(x)}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}}}P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)\\ &+S\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right).\end{array}$
Suppose that $S=V^{T}V$ for some polynomial matrix $V$. Then, we know that there exist polynomial matrices $V^{(1)},V^{(2)}$ with $2\deg(V^{(1)})\leq\deg(S)$, $2\deg(V^{(2)})+2\leq\deg(S)$ such that
$V\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)=\left(\frac{1}{\sqrt{1+\|x\|_{2}^{2}}}\right)^{\deg(V)}(V^{(1)}+\sqrt{1+\|x\|_{2}^{2}}V^{(2)}).$
Then, it holds that
$\begin{array}[]{ll}S\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)=&\frac{(V^{(1)})^{T}V^{(1)}+(1+\|x\|_{2}^{2})(V^{(2)})^{T}V^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{\deg(S)}}\\ &+\frac{(V^{(2)})^{T}V^{(1)}+(V^{(1)})^{T}V^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{\deg(S)-1}}.\end{array}$
Similarly, there exist polynomial matrices $P_{i}^{(1)},P_{i}^{(2)}$ with $\deg(P_{i}^{(1)})\leq\deg(P_{i})$, $\deg(P_{i}^{(2)})+1\leq\deg(P_{i})$ such that
$P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)=\left(\frac{1}{\sqrt{1+\|x\|_{2}^{2}}}\right)^{\deg(P_{i})}(P_{i}^{(1)}+\sqrt{1+\|x\|_{2}^{2}}P_{i}^{(2)}),$
Then, we have that
$\begin{array}[]{l}\sum\limits_{i=1}^{t}P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)^{T}\frac{G(x)}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}}}P_{i}\left(\frac{(1,x)}{\sqrt{1+\|x\|_{2}^{2}}}\right)\\ \quad\quad\quad\quad\quad\quad=\sum_{i=1}^{t}\frac{(P_{i}^{(1)})^{T}G(x)P_{i}^{(1)}+(1+\|x\|_{2}^{2})(P_{i}^{(2)})^{T}G(x)P_{i}^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}+2\deg(P_{i})}}\\ \quad\quad\quad\quad\quad\quad\,+\frac{(P_{i}^{(2)})^{T}G(x)P_{i}^{(1)}+(P_{i}^{(1)})^{T}G(x)P_{i}^{(2)}}{\sqrt{1+\|x\|_{2}^{2}}^{d_{0}+d_{G}+2\deg(P_{i})-1}}.\\ \end{array}$
The identity (5.10) implies that for $k$ satisfying (5.9), we have
$(1+\|x\|_{2}^{2})^{k}F=Q+\sqrt{1+\|x\|_{2}^{2}}H(x),$
for some $Q\in\mbox{QM}[G]^{\ell}$ with $\deg(Q)\leq 2k+d$ and symmetric polynomial matrix $H$. Since $\sqrt{1+\|x\|_{2}^{2}}$ is not a rational function, we have that $(1+\|x\|_{2}^{2})^{k}F=Q$ and thus $(1+\|x\|_{2}^{2})^{k}F\in\mbox{QM}[G]^{\ell}_{2k+d}$. ∎∎
Corollary 5.4.
Let $K$ be as in (5.1) with $m\geq 2$. Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$. If $F\succeq 0$ on $K$, then for $\varepsilon>0$, we have
$(1+\|x\|_{2}^{2})^{k}(F+\varepsilon\cdot(1+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell})\in\mbox{QM}[G]_{k}^{\ell},$
if $k\geq C\cdot\varepsilon^{-7\eta-3}.$ Here, $\eta$ is the Lojasiewicz exponent as in (5.8) and $C>0$ is a constant depending on $F$, $G$.
Proof.
Note that $F\succeq 0$ on $K$ and
$d+2\geq\deg(F+\varepsilon\cdot(1+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil})=2\lceil\frac{d+1}{2}\rceil>d.$
Hence, we know that for $\tilde{x}\in\widetilde{K}$, the homogenization of $F+\varepsilon\cdot(1+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}$ satisfies
$\begin{array}[]{ll}x_{0}^{2\lceil\frac{d+1}{2}\rceil-d}\widetilde{F}+\varepsilon\cdot(x_{0}^{2}+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}&\succeq\varepsilon\cdot(x_{0}^{2}+\|x\|_{2}^{2})^{\lceil\frac{d+1}{2}\rceil}I_{\ell}\\ &\succeq\varepsilon\cdot I_{\ell}.\\ \end{array}$
The conclusion follows from Theorem 5.3. ∎∎
5.3. Convergence rate of matrix SOS relaxations
Consider the polynomial matrix optimization
(5.11) $\left\{\begin{array}[]{rl}\min&f(x)\\ \mathit{s.t.}&G(x)\succeq 0,\\ \end{array}\right.$
where $f(x)\in\mathbb{R}[x]$, $G(x)$ is an $m\times m$ symmetric polynomial matrix. Denote by $f_{\min}$ the optimal value of (5.11).
A standard approach for solving (5.11) globally is the sum-of-squares hierarchy introduced in [15, 25, 43], consisting of a sequence of semidefinite programming relaxations. Recall that for a degree $k$, the $k$th degree truncation of $\mbox{QM}[G]$ is
$\mbox{QM}[G]_{k}:=\left\{\begin{array}[]{l|l}\sigma+\sum_{i=1}^{t}v_{i}^{T}Gv_{i}&\begin{array}[]{l}\sigma\in\Sigma[x],~{}v_{i}\in\mathbb{R}[x]^{m},\\ \deg(\sigma)\leq k,~{}\deg(v_{i}^{T}Gv_{i})\leq k\end{array}\end{array}\right\}.$
Checking whether a polynomial belongs to the the quadratic module $\mbox{QM}[G]_{k}$ can be done by solving a semidefinite program (see [43]). For an order $k$, the $k$th order SOS relaxation of (5.11) is
(5.12) $\left\{\begin{array}[]{cl}\max&\gamma\\ \text{ s.t. }&f-\gamma\in\mathrm{QM}[G]_{2k}.\end{array}\right.$
Let $f_{k}$ denote the optimal value of (5.12). It holds the monotonicity relation
$f_{k}\leq f_{k+1}\leq\cdots\leq f_{\min}.$
The hierarchy (5.12) is said to have asymptotic convergence if $f_{k}\rightarrow\infty$ as $k\rightarrow\infty$ and it has finite convergence if $f_{k}=f_{\min}$ for all $k$ big enough. Under the archimedeanness, it was shown in [15, 25, 43] that the hierarchy (5.12) has asymptotic convergence. Recently, Huang and Nie [22] proved that this hierarchy also has finite convergence if in addition the nondegeneracy condition, strict complementarity condition and second order sufficient condition hold at every global minimizer of (5.11). In the following, we provide a polynomial bound on the convergence rate of the hierarchy (5.12).
Theorem 5.5.
Suppose $f\in\mathbb{R}[x]$ with $d=\deg(f)$, $m\geq 2$, and $1-\|x\|_{2}^{2}\in\mbox{QM}[G]$. Then we have that
(5.13) $|f_{k}-f_{\min}|\leq 3\cdot(C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta})^{\frac{1}{7\eta+3}}\|f\|_{B}\left(\frac{1}{k}\right)^{\frac{1}{7\eta+3}},$
where $\kappa$, $\eta$ are respectively Lojasiewicz constant and exponent as in (5.3) and $C>0$ is a constant independent of $f$, $G$.
Proof.
Since $f_{\min}\leq\|f\|_{B}$ (see [13]), we have that for $0<\varepsilon\leq\|f\|_{B}$,
$\|f-f_{\min}+\varepsilon\|_{B}\leq\|f\|_{B}+|f_{\min}|+\varepsilon\leq 3\|f\|_{B}.$
Since $f-f_{\min}+\varepsilon\geq\varepsilon$ on $K$, it follows from Theorem 1.1 that $f-f_{\min}+\varepsilon\in\mbox{QM}[G]_{k}$ for
$k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{3\|f\|_{B}}{\varepsilon}\right)^{7\eta+3}.\$
[…]
6. Conclusions and discussions
Under the archimedeanness, we prove a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensatz. For the special case that the constraining polynomial matrix is diagonal, it improves the exponential bound shown in [14]. As a byproduct, we also obtain a polynomial bound on the convergence rate of matrix SOS relaxations, which settles an open question posed by Dinh and Pham [10]. When the constraining set is unbounded, a similar bound is provided for matrix Putinar–Vasilescu’s Positivstellensatz.
Theorem 1.1 uses the Lojasiewicz inequality (5.3) for equivalent scalar polynomial inequalities of $K$. There also exists the Lojasiewicz inequality for polynomial matrices (see [9, 10]), whose Lojasiewicz exponent is typically smaller than the scalar one as in (5.3). An interesting future work is to give a proof based on the Lojasiewicz inequality for $G(x)$ directly and it may provide a better bound. On the another hand, the polynomial matrix $F(x)\succeq 0$ on the set $K$ described by $G(x)\succeq 0$ is equivalent to that $F(x)\succeq 0$ on the quantifier set
$\{x\in\mathbb{R}^{n}\mid y^{T}G(x)y\geq 0,~{}\forall~{}\|y\|_{2}=1\}.$
It would also be interesting to study the complexity of Putinar type Positivstellensatze with universal quantifiers for the above reformulation. We refer to [18] for the recent work on Positivstellensatze with universal quantifiers.
When $K$ is given by scalar polynomial inequalities, it was shown in [2, 3] that the Lojasiewicz exponent $\eta$ as in (3.5) is equal to $1$ if the standard linear independence constraint qualification holds at every $x\in K$. Then, we can get a representation with order $\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{10}$. An analogous of the linear independence constraint qualification in nonlinear semidefinite programming is the nondegeneracy condition [47, 51]. It is also interesting to investigate if there exist similar results for polynomial matrix inequalities under the nondegeneracy condition.
Acknowledgements: The author would like to thank Prof. Konrad Schmüdgen for helpful discussions on Proposition 4.1, Prof. Tien-Son Pham for fruitful suggestions and comments which help improve this paper.